# X-Boundary: Establishing Exact Safety Boundary to Shield LLMs from Multi-Turn Jailbreaks without Compromising Usability

### v2

@@ 6 Conclusion
In this paper, we comprehensively compare existing defense methods in multi-turn attack scenarios and reveal their shortcomings in balancing the robustness of defense and LLM usability. We analyze this issue from the perspective of LLMs’ feature space, and conclude that previous methods fail to learn a precise boundary that distinguishes safe and harmful representations without an explicit formulation. To address this issue, we propose the X-Boundary to push harmful representations away from safe representations through explicit loss functions and obtain a clear distinction boundary. Such distinction boundary enables the consequential removal of harmful representations without disrupting safe ones, thereby achieving a balance between robustness against multi-turn jailbreaks and LLM usability. We think that X-Boundary can offer more efficient and fine-grained defense for LLMs, complementing existing safety alignment techniques and ultimately improving the deployment of robust AI systems in real-world applications.
Impact Statement
@@
This work aims to advance the field of large language models (LLMs) safety alignment by proposing X-Boundary, a method that maintains state-of-the-art performance in multi-turn jailbreak attack defenses while effectively mitigating the over-safety problem. We do not consider that this method will directly lead to severe negative consequences for societal development. However, we must be aware that malicious actors could exploit various approaches to induce LLMs to generate misleading or harmful content. Therefore, we expect that future research will focus on enhancing content moderation mechanisms and setting up ethical usage protocols to effectively reduce potential risks. From a constructive perspective, this method can significantly enhance the reliability, safety, and usability of LLMs.
Appendix A Additional Results
A.1 The compliance rates in HarmBench
We evaluate the basic safety level of the LLMs using HarmBench (Mazeika et al., 2024). Table 5 shows both existing defense methods and X-Boundary can effectively reduce the compliance rates to harmful requests in HarmBench without using attack methods.
@@ A.1 The compliance rates in HarmBench
Table 5: The compliance rates to harmful requests in HarmBench after using existing defense methods and X-Boundary.
Methods Llama-3- 8B-Instruct Qwen2.5- 7B-Chat Mistral-7B- Instruct-v0.2 Qwen2.5- 14B-Chat
Vanilla 11.67 26.25 56.67 15.83
SFT 1.25 5.42 7.08 1.25
DPO 0.83 2.92 13.33 5.83
GA 5.00 3.75 9.17 /
CB 1.67 4.58 13.33 4.58
X-Boundary 1.25 4.58 13.33 2.92
A.2 The effect of defense methods on the LLMs’ reasoning ability
@@
A.3 The Trade-Off between Robustness and General Capability
@@
Figure 8: The trade-off between ASR of multi-turn jailbreak and general capability on Llama-3-8B-Instruct. The data points were collected by sampling and evaluating at every 100 training steps.
A.4 Ablation Study on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2
Through analyzing the results of ablation experiments in Table 7 and Table 8, we can obtain conclusions consistent with that in Section 4.4.
A.5 Effects of the Size of Boundary-Safe Data
@@
A.6 Effects of Separate Loss and Boundary-Safe Data on the Representation Distribution
@@
Figure 10: The impact of boundary-safe data size on ASR and over-refusal rate without and with separate loss.
@@
A.7 Details about Representation Visualization
@@
Appendix B Experimental Details
B.1 Construction of Multi-Turn Defense Dataset
We construct a multi-turn defense dataset based on SafeMTData. SafeMTData is derived from the circuit breaker training dataset, and carefully filtered to prevent data contamination with Harmbench. It includes harmful multi-turn queries generated by ActorAttack (Ren et al., 2024b), along with refusal responses to reject the harmful queries. To curate the harmful responses for DPO, GA, and CB, we use harmful multi-turn queries in SafeMTData to attack deepseek-chat (Liu et al., 2024a) and filter the harmful response using HarmBench classifier (Mazeika et al., 2024).
For SFT and DPO, following Ren et al. (2024b), we maintain a 1:2 ratio between the multi-turn defense data and instruction data, e.g., UltraChat (Ding et al., 2023). For CB, we add the filtered harmful responses and their corresponding single-turn queries into the CB dataset. The other data settings remain consistent with Zou et al. (2024). For GA, we follow (Zhang et al., 2024b) and use unlearning data, instruction data, and refusal data in a ratio of 5:5:1.
B.2 Training Details of Baselines
@@
Following the experimental setting of Zhang et al. (2024b), we set the batch size to 11 with accumulation step 1, where the ratio of the three types of data in a batch is 5:5:1. We use the AdamW optimizer with a learning rate of $2\times 10^{-5}$ and set the maximum epoch as 3. For Qwen2.5-7B-Chat and Llama-3-8B-Instruct, the coefficients of safe responses loss $\mathcal{L}_{s}$, general performance loss $\mathcal{L}_{g}$, and unlearning loss $\mathcal{L}_{h}$ are set to 0.5, 1.0, 0.3. For Mistral-7B-Instruct-v0.2, the loss coefficients are set to 0.25, 1.0, and 0.05, respectively. All training processes use Low-Rank Adaptation (LoRA) for parameter fine-tuning. For Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2, we set the rank $r$, scaling factor $\alpha$, and dropout rate to 16, 16, 0.05. For Qwen2.5-7B-Chat, we conducted a grid search over the LoRA hyperparameters with $r\in\{8,16,32\}$ and $\alpha\in\{16,32,64\}$. We end up selecting $r=8$, $\alpha=64$, and a dropout rate of $0.05$. We linearly decay the learning rate and select the checkpoint after 1 epoch for evaluation. Training a Mistral-7B-Instruct-v0.2 model on a single A100 80GB GPU takes approximately 1 hour.
@@
We follow (Zou et al., 2024) to use LoRA for fine-tuning and set the rank $r$ as 16 on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2, 32 on Qwen2.5-7B-Chat and Qwen2.5-14B-Chat. We gather the feature representations from layers 10, 20, 30, and 40 to calculate circuit-breaking loss and inset LoRA adapter into all linear layers from 0 through 40. The loss coefficients are dynamically adjusted. The coefficients of circuit-breaking loss and retain loss are $c_{s}=\alpha(1-\frac{t}{\beta})$ and $c_{r}=\alpha\frac{t}{\beta}$, respectively. We set $\alpha$ as 5 on Mistral-7B-Instruct-v0.2 and 10 on other LLMs, $\beta$ as 300 on Mistral-7B-Instruct-v0.2 and Llama-3-8B-Instruct, 600 on Qwen2.5-7B-Chat, and 1200 on Qwen2.5-14B-Chat. Qwen2.5-14B-Chat is trained on for 360 steps with a batch size of 8 on 4 A100 GPUs, while other LLMs is trained on for 180 steps with a batch size of 16 on 1 A100 GPU.
B.3 Training Details of X-Boundary
We use LoRA for fine-tuning and set the rank $r$ as 16 on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2, 32 on Qwen2.5-7B-Chat and Qwen2.5-14B-Chat. We set dynamic loss coefficients following (Zou et al., 2024), where $c_{r}=\alpha\frac{t}{\beta}$ and $c_{e}=c_{s}=\alpha(1-\frac{t}{\beta})$. $\alpha,\beta$, and the target layers for calculating erase loss keep consistent with hyperparameters specified in Appendix B.2. We conduct a grid search on the size of boundary-safe data in a valid set in the range of [0,500], with a step of 50, selecting the size for Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.2, Qwen2.5-7B-Chat, and Qwen2.5-14B-Chat is 500, 200, 100, and 50, respectively. Qwen2.5-14B-Chat is trained for 260 steps with a batch size of 8 on 4 A100 GPUs, while other LLMs are trained for 180 steps with a batch size of 16 on 1 A100 GPU.
B.4 Evaluations
Datasets.

### v3

@@ 5 Conclusion
In this paper, we comprehensively compare existing jailbreak defense methods and reveal the trade-off between the robustness of defense and LLM usability. We analyze this issue from the perspective of LLMs’ feature space, and conclude that previous methods fail to learn a precise boundary that distinguishes safe and harmful representations without an explicit formulation. To address this issue, we propose X-Boundary to push harmful representations away from safe representations through explicit loss functions and obtain a clear distinction boundary. Such distinction boundary enables the consequential removal of harmful representations without disrupting safe ones, thereby achieving a balance between robustness against jailbreaks and LLM usability. We think that X-Boundary can offer a more efficient and fine-grained defense for LLMs, improving the deployment of robust AI systems in real-world applications.
Limitations
This paper has several limitations. First, although we analyze the underlying causes of the trade-off between defense robustness and LLM usability and propose a post-training method to achieve a mutually beneficial outcome, we have not yet thoroughly investigated how to fundamentally resolve this issue during the pre-training stage, as the pre-training processes of these LLMs are closed-source. Second, due to its reliance on representation-level intervention, X-Boundary is not applicable to black-box models, thereby restricting its use in some practical settings.
Ethical considerations
@@
This work aims to advance the field of large language models (LLMs) safety alignment by proposing X-Boundary, a method that maintains state-of-the-art performance in multi-turn jailbreak attack defenses while effectively mitigating the over-safety problem. All the training data and reproduced defense methods we used are open-source and consistent with their intended use, with proper citations to their original sources. We do not consider that this method will directly lead to severe negative consequences for societal development. However, we must be aware that malicious actors could exploit various approaches to induce LLMs to generate misleading or harmful content. Besides, training data containing some harmful or offensive questions and answers pose a risk of malicious use and potential harm. Therefore, we expect that future research will focus on enhancing content moderation mechanisms and setting up ethical usage protocols to effectively reduce potential risks.
Acknowledgements
This work is supported by Shanghai Artificial Intelligence Laboratory. And we would like to express our gratitude to our collaborators for their efforts.
Appendix A Related Work
Jailbreak attacks. Jailbreak attacks aim to bypass the safety mechanisms of large language models (LLMs), prompting them to generate harmful or policy-violating content defense_survey. These attacks can be broadly categorized into single-turn and multi-turn scenarios based on their interaction structure with the model backdoor_multi; scaleAI_multi. One representative method is GCG zou2023GCG, which formulates jailbreak as an optimization problem and employs genetic algorithms to automatically evolve effective attack prompts. AutoDAN autodan automates the generation of adversarial prompts through a dynamic prompt-injection framework and achieves high attack success rates with minimal human intervention. Unlike single-turn jailbreaks, multi-turn jailbreaks exploit flexible multi-turn dialogues to bypass the safeguards of LLMs bju_multi; imposterAI_multi; red_queen, making them challenging to detect and defend against. For example, cosafe_multi, bju_multi and imposterAI_multi generate multi-turn jailbreak queries by breaking down the original malicious query into multiple less harmful sub-questions. actor_attack; coa and crescendo dynamically adjust the attack query based on the contextual feedback from victim LLMs, gradually steering benign initial queries toward more harmful topics throughout the conversation.
Defenses for LLMs. Although defense methods for multi-turn jailbreak attacks are less explored in the literature, some existing approaches have proven effective against various single-turn attacks and have the potential to be adapted for multi-turn scenarios. These defense methods can be classified into the following categories: training LLMs to refuse harmful queries (bai2022training; dpo; ouyang2022training; decoupled_sft), training LLMs to prioritize safe instructions (lu2024sofa; wallace2024instruction; zhang2023defending), unlearning and editing harmful knowledge  (eraser; safe_unlearning; ren2024identifying; cq), prompt engineering (xie2023defending; zheng2024prompt), and implementing input and output guardrails (inan2023llama; dubey2024llama) such as jailbreak detection (hu2024gradient; jain2023baseline) input perturbation (cao2023defending; robey2023smoothllm; liu2024protecting). Several studies  (li2024wmdp; circuit_breaker; zou2023representation; qian2024towards; zhang2024better) also propose defense methods from the perspective of representation engineering, inspiring us to optimize LLMs in the representation space to strike a balance between defense robustness and LLM usability.
Decline in usability caused by defense methods. We assess the impact of defense methods on usability from two aspects: general capability degradation and over-refusal. General capability degradation, commonly known as the “alignment tax” alignment_tax phenomenon, has garnered widespread attention and has been extensively discussed in technical reports on LLMs dubey2024llama; inan2023llama; actor_attack; li2024wmdp; vlsbench. Over-refusal refers to the unreasonable rejection of safe queries by LLMs varshney2023art; zhao2024towards; zou2023representation; arditi2024refusal; cao2024nothing. bianchi2023safety discover that excessive safety-tuning makes LLMs refuse entirely safe prompts if they superficially resemble unsafe ones. xstest, oktest, orbench, and phtest employ linguistic techniques or automatic pipelines to generate seemingly unsafe prompts for evaluating LLMs’ over-refusal behavior. Previous studies have explored several approaches to mitigate over-refusal. For example, oktest applied contrastive decoding by inferencing twice on the same query with and without the system prompt. wang2024surgical extract and ablate a false refusal vector to reduce over-refusal rate. In this paper, we evaluate the performance of these methods and compare them with X-Boundary.
Appendix B The Optimization Process of X-Boundary
The optimization process of X-Boundary is shown as Algorithm 1.
Algorithm 1 The optimization process of X-Boundary
1:for $t=1$ to $T$ do
3:  Sample $\{(x_{i}^{b},x_{i}^{r})\}_{i=1}^{n}\sim\mathcal{D}_{b}$
4:  $c_{r}=\alpha\frac{t}{\beta}$, $c_{e}=c_{s}=\alpha(1-\frac{t}{\beta})$
5:  $\mathcal{L}_{\text{r}}=\frac{1}{n}\sum_{i=1}^{n}\left\|\mathcal{R}_{\mathcal{M}_{\theta}}\left(x_{i}\right)-\mathcal{R}_{\mathcal{M}_{\texttt{ref}}}\left(x_{i}\right)\right\|_{2}$
8:  $\mathcal{L}=c_{r}\mathcal{L}_{r}+c_{e}\mathcal{L}_{e}+c_{s}\mathcal{L}_{s}$
9:  Update parameters $\theta$ to minimize $\mathcal{L}$
10:end for
Appendix C Additional Results
C.1 Evaluation of Existing Over-Refusal Mitigation Methods
To further investigate the trade-off issue, we implement three over-refusal mitigation methods: system prompt (SP) oktest, Self-Contrastive Decoding (Self-CD) oktest, and vector ablation (VA) wang2024surgical. Table 3 shows that these methods are effective on the vanilla model (Qwen2.5-7B-Instruct) and do not lead to a significant increase in ASR. However, as shown in Table 4, their impact on reducing ORR is less noticeable in models fine-tuned with defense methods, and they substantially weaken the defense effectiveness. Furthermore, both Self-CD and VA depend on refusal vectors or refusal tokens, which are ineffective for methods like CB that do not use a fixed refusal template.
Methods Attack Success Rate (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
DirectRequest ActorAttack XSTest OKTest OR-Bench PHTest MMLU GSM8K HumanEval
Qwen2.5-7B-Instruct 26.25 76.00 6.00 19.33 1.67 5.67 74.26 80.67 81.71
+SP 26.67 78.50 2.80 9.33 1.67 3.67 74.30 80.97 81.10
+Sefl-CD 28.33 78.00 2.80 9.33 1.00 4.33 74.21 80.52 82.93
+VA 27.92 75.50 4.20 11.00 1.33 3.00 74.58 80.36 81.71
Table 3: Performance of existing over-refusal mitigation methods on Qwen2.5-7B-Instruct.
Methods Attack Success Rate (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
DirectRequest ActorAttack XSTest OKTest OR-Bench PHTest MMLU GSM8K HumanEval
Qwen2.5-7B-Instruct 26.25 76.00 6.00 19.33 1.67 5.67 74.26 80.67 81.71
+SFT 5.42 21.00 46.00 57.67 29.33 53.67 74.30 76.42 77.44
+SFT+SP 6.25 41.00 37.20 47.00 26.00 44.00 74.17 75.51 78.66
+SFT+Sefl-CD 6.00 28.50 44.80 52.67 28.33 54.00 73.63 77.94 79.27
+SFT+VA 8.75 43.50 23.60 41.33 23.67 40.00 74.58 77.94 78.66
+CB 1.67 15.50 20.60 26.00 34.00 43.67 74.21 80.36 81.10
+CB+SP 2.92 27.00 20.20 27.33 35.67 42.00 74.21 80.43 80.49
+CB+Sefl-CD 4.58 26.50 24.80 25.00 37.33 46.33 74.30 80.52 79.88
+CB+VA 2.08 20.50 19.20 24.00 33.67 41.33 73.67 80.43 80.49
X-Boundary 1.25 17.50 10.40 16.67 5.33 15.00 74.17 80.52 81.10
Table 4: Performance of existing over-refusal mitigation methods on Qwen2.5-7B-Instruct fine-tuned with defense methods.
C.2 Performance on Qwen2.5-14B-Instruct
Table 5 shows that X-Boundary also achieves SOTA defense and the lowest ORR on Qwen2.5-14B-Instruct.
Methods Attack Success Rate (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
DirectRequest ActorAttack RedQueen Crescendo XSTest OKTest OR-Bench PHTest MMLU GSM8K HumanEval
Vanilla 15.83 71.50 63.50 36.00 4.00 10.00 1.33 4.00 80.06 82.49 79.88
SFT 7.08 52.00 10.00 16.00 43.60 51.33 31.33 62.67 79.58 82.18 81.71
DPO 8.33 54.50 45.00 32.00 6.40 14.00 2.67 8.67 78.58 83.32 81.10
CB 3.33 23.50 4.50 8.00 43.60 51.33 32.00 64.33 79.64 82.56 82.93
X-Boundary 2.91 25.00 5.00 12.00 5.20 13.67 4.00 8.33 79.52 82.18 81.10
Table 5: Comparison of existing defense methods and X-Boundary on Qwen2.5-14B-Instruct.
C.3 Defense Performance Against Single-Turn Jailbreak Attacks
We evaluate the robustness of X-Boundary and baseline methods against seven single-turn jailbreak attacks, i.e., GCG zou2023GCG, PAIR chao2023PAIR, PAP zeng2024PAP, AutoDAN autodan, Obfuscation zhang2024obfuscation, Spliting kang2023spliting, and Multilingual yong2023translation. Table 6 shows X-Boundary can effectively reduce the ASR of these attacks.
Methods DirectRequest GCG PAIR PAP AutoDAN Obfuscation Splitting Multilingual
Vanilla 11.67 31.00 18.00 15.00 4.50 12.00 15.00 3.00
SFT 1.25 6.50 13.50 1.50 0.50 2.00 7.00 0.00
DPO 0.83 8.50 11.00 3.00 0.00 4.00 1.00 0.00
GA 5.00 18.00 11.50 3.50 1.50 9.50 7.00 1.00
CB 1.67 2.00 12.00 1.00 0.00 0.00 2.00 0.00
X-Boundary 1.25 1.50 10.00 1.00 0.00 0.50 3.00 0.00
@@ C.3 Defense Performance Against Single-Turn Jailbreak Attacks
Table 6: The ASR of seven single-turn jailbreak attacks after using existing defense methods and X-Boundary.
C.4 The Effect of Defense Methods on the LLMs’ Reasoning Ability
@@
C.5 The Trade-Off between Robustness and General Capability
@@
The trade-off between ASR of multi-turn jailbreak and general capability on Llama-3-8B-Instruct. The data points were collected by sampling and evaluating at every 100 training steps.
C.6 Ablation Studies on Three Loss Terms
We conduct ablation studies on three loss terms on Llama-3-8B-Instruct. Table 8 indicates that three losses all contribute significantly to performance. Specifically, the erase loss $\mathcal{L}_{\texttt{e}}$ primarily reduces the ASR, while the retain loss $\mathcal{L}_{\texttt{r}}$ maintains general capabilities without significant degradation and prevents a substantial increase in the ORR. Additionally, the separate loss $\mathcal{L}_{\texttt{s}}$ further preserves general capabilities, reduces the ORR, and ensures the overall usability of the model.
Jailbreak ASR (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
DirectRequest ActorAttack XSTest OKTest OR-Bench PHTest MMLU GSM8K HumanEval
Vanilla 11.67 58.50 6.80 9.00 8.00 13.67 68.30 79.08 59.18
w/o $\mathcal{L}_{\texttt{e}}$ 12.50 57.00 5.60 8.33 6.67 14.00 68.30 80.21 59.76
w/o $\mathcal{L}_{\texttt{r}}$ 0.00 0.00 100.00 100.00 100.00 100.00 68.30 77.86 57.32
w/o $\mathcal{L}_{\texttt{s}}$ 1.67 16.50 23.60 27.67 36.00 52.00 67.67 78.47 59.76
X-Boundary 1.25 16.50 8.40 14.00 8.00 28.67 67.94 78.70 59.76
Table 8: Evaluation results comparing different model settings.
C.7 Sensitivity Analysis on Hyper-Parameters
We analyze the sensitivity analysis on hyper-parameters $\alpha$ and $\beta$, where $\mathcal{L}=c_{r}\mathcal{L}_{r}+c_{e}\mathcal{L}_{e}+c_{s}\mathcal{L}_{s}$, $c_{e}=c_{s}=\alpha(1-\frac{t}{\beta}$ and $c_{r}=\alpha\frac{t}{\beta}$. Specifically, we vary $\alpha\in\{5,10,15,20\}$ and $\beta\in\{200,250,300,350\}$. Fig. 12 shows that X-Boundary is relatively insensitive to $\alpha$. As the hyper-parameter $\beta$ increases, i.e., meaning the coefficients of the erase loss $\mathcal{L}_{e}$ and separate loss $\mathcal{L}_{s}$ are scaled up while the coefficient of the retain loss $\mathcal{L}_{r}$ are scaled down, the ASR tends to decrease, while the ORR tends to rise.
Sensitive analysis on hyper-parameters $\alpha$ and $\beta$.
C.8 Ablation Studies on Three Models
Through analyzing the results of ablation experiments in Table 9, Table 10 and Table 11, we can obtain conclusions consistent with that in Section 4.4.
C.9 Effects of the Size of Boundary-Safe Data
@@
C.10 Effects of Separate Loss and Boundary-Safe Data
@@
Models A B C D Single & Multi-Turn ASR (%) $\downarrow$ Over-Refusal Rate (%) $\downarrow$ General Capability (%) $\uparrow$
DirectRequest ActorAttack XSTest OKTest OR-Bench PHTest MMLU GSM8K HumanEval
Vanilla 15.83 71.50 4.00 10.00 1.33 4.00 80.06 82.49 79.88
(a) $\checkmark$ 4.17 56.50 6.00 9.00 4.33 7.00 79.64 82.95 81.10
(b) $\checkmark$ $\checkmark$ 2.92 31.00 12.80 19.67 53.00 48.33 79.65 83.25 80.49
(c) $\checkmark$ $\checkmark$ $\checkmark$ 4.17 31.00 8.40 16.00 9.33 16.33 79.48 83.33 80.49
X-Boundary $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 2.92 25.00 5.20 13.67 4.00 8.33 79.52 82.18 81.10
Table 11: Ablation study on Qwen2.5-14B-Instruct. In this table, A represents single-turn defense data, B represents multi-turn defense data, C represents boundary-safe data, and D represents the separate loss $\mathcal{L}_{\texttt{s}}$.
The impact of boundary-safe data size on ASR and over-refusal rate without and with separate loss.
@@
C.11 Details about Representation Visualization
@@
Appendix D Experimental Details
D.1 Construction of Multi-Turn Defense Dataset
We construct a multi-turn defense dataset based on SafeMTData. SafeMTData is derived from the circuit breaker training dataset, and carefully filtered to prevent data contamination with Harmbench. It includes harmful multi-turn queries generated by ActorAttack actor_attack, along with refusal responses to reject the harmful queries. To curate the harmful responses, we use harmful multi-turn queries in SafeMTData to attack deepseek-Instruct deepseek and filter the harmful response using HarmBench classifier harmbench.
For SFT, we directly exploit SafeMTData as a multi-turn training dataset following actor_attack. For DPO, we follow red_queen to construct preference pair using curated harmful responses and refusal response in SafeMTDate as rejected and chosen data, respectively. For SFT and DPO, we follow actor_attack to maintain a 1:2 ratio between the multi-turn defense data and instruction-following data, e.g., UltraChat ultrachat. For CB, we add pairs of harmful queries from SafeMTData along with the curated harmful responses into its defense training datasets to remove harmful knowledge that could be elicited through multi-turn attacks. The other data settings remain consistent with circuit_breaker. For GA, we add harmful queries from SafeMTData along with the curated harmful responses to the unlearning dataset and follow safe_unlearning to use unlearning data, instruction-following data, and refusal data in a ratio of 5:5:1.
D.2 Training Details of Baselines
We compare X-Boundary with the following four methods:
  • Multi-Turn SFT actor_attack: fine-tuning LLMs using harmful queries as inputs and refusal answers as supervised labels directly.
  • Multi-Turn DPO dpo; red_queen: aligning LLMs using harmful queries as inputs, harmful answers as rejected responses, and refusal answers as chosen responses.
  • GA safe_unlearning; eraser: unlearning harmful knowledge by training with gradient ascent optimization methods .
  • CB circuit_breaker: remapping the representations of harmful knowledge to desired targeted representations.
@@ D.2 Training Details of Baselines
Following the experimental setting of safe_unlearning, we set the batch size to 11 with accumulation step 1, where the ratio of the three types of data in a batch is 5:5:1. We use the AdamW optimizer with a learning rate of $2\times 10^{-5}$ and set the maximum epoch as 3. For Qwen2.5-7B-Instruct and Llama-3-8B-Instruct, the coefficients of safe responses loss $\mathcal{L}_{s}$, general performance loss $\mathcal{L}_{g}$, and unlearning loss $\mathcal{L}_{h}$ are set to 0.5, 1.0, 0.3. For Mistral-7B-Instruct-v0.2, the loss coefficients are set to 0.25, 1.0, and 0.05, respectively. All training processes use Low-Rank Adaptation (LoRA) for parameter fine-tuning. For Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2, we set the rank $r$, scaling factor $\alpha$, and dropout rate to 16, 16, 0.05. For Qwen2.5-7B-Instruct, we conducted a grid search over the LoRA hyperparameters with $r\in\{8,16,32\}$ and $\alpha\in\{16,32,64\}$. We end up selecting $r=8$, $\alpha=64$, and a dropout rate of $0.05$. We linearly decay the learning rate and select the checkpoint after 1 epoch for evaluation. Training a Mistral-7B-Instruct-v0.2 model on a single A100 80GB GPU takes approximately 1 hour.
@@
We follow circuit_breaker to use LoRA for fine-tuning and set the rank $r$ as 16 on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2, 32 on Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct. We gather the feature representations from layers 10, 20, 30, and 40 to calculate circuit-breaking loss and inset LoRA adapter into all linear layers from 0 through 40. The loss coefficients are dynamically adjusted. The coefficients of circuit-breaking loss and retain loss are $c_{s}=\alpha(1-\frac{t}{\beta})$ and $c_{r}=\alpha\frac{t}{\beta}$, respectively. We set $\alpha$ as 5 on Mistral-7B-Instruct-v0.2 and 10 on other LLMs, $\beta$ as 300 on Mistral-7B-Instruct-v0.2 and Llama-3-8B-Instruct, 600 on Qwen2.5-7B-Instruct, and 1200 on Qwen2.5-14B-Instruct. Qwen2.5-14B-Instruct is trained on for 360 steps with a batch size of 8 on 4 A100 GPUs, while other LLMs is trained on for 180 steps with a batch size of 16 on 1 A100 GPU.
D.3 Training Details of X-Boundary
We use LoRA for fine-tuning and set the rank $r$ as 16 on Llama-3-8B-Instruct and Mistral-7B-Instruct-v0.2, 32 on Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct. We set dynamic loss coefficients following circuit_breaker, where $c_{r}=\alpha\frac{t}{\beta}$ and $c_{e}=c_{s}=\alpha(1-\frac{t}{\beta})$. $\alpha,\beta$, and the target layers for calculating erase loss keep consistent with hyperparameters specified in Appendix D.2. We conduct a grid search on the size of boundary-safe data in a valid set in the range of [0,500], with a step of 50, selecting the size for Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.2, Qwen2.5-7B-Instruct, and Qwen2.5-14B-Instruct is 500, 200, 100, and 50, respectively. The training deploys the AdamW optimizer with a fixed learning rate of 1e-4. Qwen2.5-14B-Instruct is trained for 260 steps with a batch size of 8 on 4 A100 GPUs, while other LLMs are trained for 180 steps with a batch size of 16 on 1 A100 GPU.
D.4 Comparison of Computational Resource Consumption
Table 12 presents a comparison of computational resource consumption with existing algorithms. The training time and VRAM Usage are tested on an A100 GPU.
Method Training Time (h) $\downarrow$ VRAM Usage (GB) $\downarrow$ Average ASR (%) $\downarrow$ Average ORR (%) $\downarrow$
SFT 0.66 23.79 8.25 37.22
DPO 1.93 67.65 9.83 26.67
GA 1.14 75.33 14.17 18.62
CB 0.49 46.65 14.17 34.82
X-Boundary 0.45 46.75 6.67 14.77
Table 12: The comparison of computational resource consumption.
D.5 Evaluations
Datasets